# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ Abstract
Gauge symmetries and Higgs mechanisms are key features of theories describing high-energy particle physics and collective phenomena in statistical and condensed-matter physics. In this review we address the collective behavior of systems of multicomponent scalar fields interacting with gauge fields, which can be already present in the underlying microscopic system or emerge only at criticality. The interplay between local gauge and global symmetries determines the phase diagram, the nature of the Higgs phases, and the nature of phase transitions between the high-temperature disordered and the low-temperature Higgs phases. However, further crucial features determine the universal properties of the critical behaviors at continuous transitions. Specifically, their nature also depends on the role played by the gauge modes at criticality: effective (Abelian or non-Abelian) gauge Higgs field theories emerge if gauge modes develop critical correlations, while more standard critical behaviors occur in the case they remain short ranged at the transition. In the latter case, gauge fields only prevent non-gauge invariant correlators from becoming critical. This review covers the recent progress made in the study of Higgs systems with Abelian and non-Abelian gauge fields. We mainly focus on three-dimensional systems, and briefly discuss two-dimensional systems, whose equilibrium statistical properties are described by a classical partition function. However, by using the quantum-to-classical mapping, the results on the critical behavior for classical systems in $D=d+1$ dimensions can be extended to quantum transitions in $d$ dimensions.
@@
Gauge theories with scalar fields, Abelian gauge Higgs theory, non-Abelian gauge Higgs theory, Higgs phases, Phase transitions, Critical behaviors, Topological transitions
@@ 1.1 Introduction
Many collective phenomena in condensed-matter physics are described by effective Abelian Higgs (AH) theories, which describes a $d$-dimensional system of degenerate $N$-component scalar fields minimally coupled with an Abelian U(1) gauge field. We mention transitions in superconductors, see, e.g., Refs. [18, 19], and in quantum SU($N$) antiferromagnets, see, e.g., Refs. [27, 28, 29, 30, 31, 32, 20], and some unconventional quantum transitions characterized by the so-called deconfined quantum criticality, see, e.g., Refs. [33, 34, 35, 36, 37, 38, 39, 40]. The phase structure and universal features of AH models have been extensively studied, see, e.g., Refs. [18, 19, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 35, 76, 77, 78, 30, 79, 31, 80, 32, 81, 82, 83, 84, 20, 85, 86, 87, 88, 89, 90, 91, 21, 92, 93], paying particular attention to the role of the gauge fields and of the related topological features, like monopoles and Berry phases, which cannot be captured by effective Landau-Ginzburg-Wilson (LGW) theories with gauge-invariant scalar order parameters [5, 94, 6, 2, 95], see, e.g., Refs. [33, 20, 21].
@@
From the theoretical point of view, the existence of continuous phase transitions that cannot be described by the standard gauge-invariant LGW paradigm is related to the nonperturbative infrared behavior of strongly coupled gauge theories, and, in particular, to the existence of infrared stable fixed points (FPs) of their renormalization-group (RG) flow. At a fundamental level, the existence of these nontrivial transitions also provides a way to define gauge field theories in a nonperturbative framework.
Over the past decade, researchers with different backgrounds, working in high-energy, statistical, and condensed-matter physics, have made important steps forward in the theoretical understanding of the nonperturbative physics of Higgs models in less than four dimensions. In light of this steady progress, a survey of the field seems to be useful for two different reasons: First, one can present the state of the art in the field; second, there is the possibility to collect and discuss in a unified fashion results that appeared in different contexts. This will also promote a more efficient communication and interdisciplinary interactions among researchers working in different areas.
To the best of our knowledge, a review on lower-dimensional, i.e., three-dimensional (3D) and two-dimensional (2D), AH and NAH theories is still missing, and we would like to fill such gap. Of course, some of the aspects we plan to cover have already been reviewed— emerging gauge fields and deconfined criticality are obvious examples, see, e.g., Refs. [20, 21]. However, the existing reviews focus on specific issues and do not report a general discussion of lower-dimensional lattice AH and NAH theories. In this review we thoroughly examine the phenomenology of gauge systems, in particular their phase diagram and the nature of the critical transitions observed when varying the gauge and global symmetries.
@@ 1.2 Plan
This review focuses on lower-dimensional AH and NAH theories, in which multicomponent scalar fields are coupled with Abelian and non-Abelian gauge fields. We discuss these models in a statistical framework [100, 2, 10, 3, 4], in which the partition function is defined as a sum over classical lattice configurations, thus providing a regularization of the functional path-integral and, possibly, paving the way for a nonperturbative definition of the corresponding statistical field theory. The quantum-to-classical mapping allows us to apply the classical results for the critical behavior (equivalently, continuum limit) in $(d+1)$ dimensions to quantum transitions of many-body systems in $d$ dimensions, see, e.g., Refs. [14, 101]. In particular, since we consider isotropic $(d+1)$-dimensional statistical models, the results can be applied to the large class of continuous quantum transitions characterized by a dynamic exponent $z=1$, i.e., to transitions in which the critical gap decays as the inverse of the spatial size [14, 101]. We mainly consider 3D classical systems, but we also briefly discuss 2D systems.
@@
The review covers the recent developments in the studies of lattice AH and NAH systems (LAH and LNAH, respectively). It discusses their phase diagram, the main features of their Higgs ordered phases [102, 103, 104, 105], where gauge correlations are gapped and nonlocal charged operators may condense, and the nature of their phase transitions, paying particular attention to transitions characterized by a nontrivial interplay of gauge and scalar modes. We will show that the standard effective LGW $\Phi^{4}$ field theory approach [94, 2, 95], which is commonly applied to critical phenomena in statistical systems without gauge symmetries, is not always able to describe transitions in gauge models. Other different scenarios indeed emerge. We pay particular attention to the comparison with the predictions of the corresponding AH and NAH field theories (AHFTs and NAHFTs, respectively), to understand when, and how, the critical behavior of lattice systems is effectively described by the field-theory model. If this is the case, the lattice model provides a way, the so-called continuum limit, to define nonperturbatively the field theory.
@@
We begin by discussing the possible types of critical behavior that emerge at continuous transitions in the presence of gauge symmetries, in Secs. 2 and 3. This requires an extension of the standard LGW paradigm, in particular when gauge modes play an active role at criticality. Secs. 4-6 focus on systems characterized by an Abelian U(1) gauge invariance, i.e., the AHFTs and the corresponding LAH models, considering models with both noncompact and compact gauge fields, which turn out to present significantly different features. Lattice models with discrete Abelian gauge symmetries are addressed in Sec. 7. Secs. 8 and 9 focus instead on field theories and lattice systems with non-Abelian gauge symmetries. Two-dimensional models, with Abelian and non-Abelian gauge symmetries, are discussed in Sec. 10. Finally, in Sec. 11 we discuss the effects of perturbations explicitly breaking the gauge symmetry, an issue which is particularly relevant for systems in which the gauge symmetry effectively emerges at the transitions but which are gauge invariant at the microscopic level. A more detailed description of the content of the review follows.
@@
    In Sec. 3, we discuss the key features of the critical behavior of statistical systems with local gauge symmetries. We argue that different types of critical behaviors may be realized. There are transitions driven by scalar fields whose nature depends on the role played by the gauge modes, in particular whether they develop or not critical correlations, and transitions driven by topological gauge modes without local order parameters. The effective description of these different critical behaviors requires different effective approaches, such as the standard gauge-invariant LGW approach when gauge fields are not critical, or the more complex gauge field theory (GFT) approach with explicit gauge fields in the presence of critical gauge correlations. This distinction is crucial to understand which statistical field theory (SFT) is actually realized by the various continuous phase transitions of LAH and LNAH models.
@@
    In Sec. 4 we focus on AHFTs, or scalar quantum electrodynamics (QED), which describe $d$-dimensional systems of degenerate $N$-component scalar fields minimally coupled with Abelian U(1) gauge fields. In particular, we consider systems with global SU($N$) and SO($N$) symmetries, depending on the symmetry of the quartic self-interactions of the scalar fields. We review the main features of their RG flow, whose stable charged FP is expected to describe the critical behavior at charged continuous transitions, when both scalar and gauge fields are critical.
@@
    In Sec. 5 we analyze lattice models corresponding to the AHFTs discussed in Sec. 4, that are characterized by the same global and local symmetries. We consider 3D LAH models in which a noncompact Abelian gauge field is coupled with an $N$-component scalar field, providing a straightforward lattice discretization of the AHFT. We review results for their phase diagrams, which present various phases, including a Higgs phase, and for their critical behaviors along the various transition lines. Different critical behaviors occur, which can be described by using the different approaches outlined in Sec. 3. In particular, the transitions separating the Coulomb and Higgs phases realize the critical continuum limit associated with the RG flow of the AHFTs for a sufficiently large number of scalar components.
@@
    In Sec. 8 we review NAHFTs, in which multiflavor scalar fields are coupled with non-Abelian gauge fields. We discuss models with SU($N_{c}$) and SO($N_{c}$) gauge symmetries, with scalar fields transforming in different representations of the gauge group, such as the fundamental and adjoint representations. We outline the known features of their RG flow, whose stable infrared FPs are expected to describe the corresponding critical behaviors at continuous phase transitions in statistical models.
@@
    In Sec. 9 we consider LNAH models, which are lattice non-Abelian gauge theories with multicomponent degenerate scalar fields. We present results for their phase diagram, paying particular attention to their low-temperature Higgs phases. We discuss the interplay between non-Abelian gauge symmetries and global symmetries, which crucially determines the properties of the various Higgs phases, and therefore the nature of the phase transitions between the disordered and the Higgs phases. We consider in particular SU($N_{c}$) gauge theories with multicomponent scalar fields in the fundamental and adjoint representations of the gauge group, two choices which lead to different Higgs phases. Some of the transitions between the disordered and the Higgs phases are continuous and can be associated with the charged FP of the corresponding NAHFT. We close this section by briefly discussing SO($N_{c}$) gauge theories with multiflavor scalar matter.
@@
    In Sec. 10 we review results for the critical continuum limit of 2D LAH and LNAH models with continuous global symmetries. They show that also the critical continuum limit of 2D LAH and LNAH models in the zero-temperature limit arises from a nontrivial interplay between gauge and global symmetries. The emerging scenarios lead us to conjecture that the zero-temperature critical behavior, and therefore the continuum limit, of 2D lattice gauge models with scalar fields belongs to the zero-temperature universality classes associated with a 2D field theory associated with one of the symmetric spaces that have the same global symmetry.
@@
    In Sec. 11 we discuss the effects of perturbations that give rise to an explicit gauge-symmetry breaking (GSB) in lattice gauge theories. This issue is discussed in LAH models with noncompact and compact gauge fields in the presence of photon-mass terms. We argue that their effect is irrelevant at the continuous transitions where gauge correlations are not critical and the critical behavior can be described by effective LGW theories with gauge-invariant order-parameter fields.
@@
    In A we report an overview of the homogeneous scaling laws expected at generic continuous phase transitions, as inferred by the RG theory of critical phenomena. We report the RG scaling relations valid in the thermodynamic limit and in the finite-size scaling (FSS) limit, at critical and multicritical points. We also briefly discuss FSS at first-order transitions.
@@
    For reference, in B we report the presently most accurate estimates of the critical exponents of some 3D O($N$)-vector universality classes, which are often referred to in this review, such as the 3D Ising, XY, and O(3) vector (Heisenberg) universality classes.
@@
    In C we present some details on the mean-field analyses of the Higgs phases of LNAH models with SU($N_{c}$) gauge symmetry and $N_{f}$ flavor in the fundamental representation.
@@
Another important issue that is not addressed by this review concerns the new features that may arise when fermionic fields are added. This topic is quite interesting, since field-theoretical models with emerging gauge symmetries in condensed-matter physics often involve fermionic excitations beside scalar modes, see, e.g., Refs. [96, 20]. We believe that this issue may provide further interesting developments, that would extend our present understanding of the phenomenology of the transitions in the presence of gauge symmetries. However, fermionic models have been much less studied so far, so we prefer to restrict our review to gauge models with scalar matter only, for which there is already a large amount of results.
In the following we report a list of the abbreviations and symbols used throughout this review (in alphabetic order):

### v3

# Three-dimensional Abelian and non-Abelian gauge Higgs theories

@@ Abstract
Gauge symmetries and Higgs mechanisms are key features of theories describing high-energy particle physics and collective phenomena in statistical and condensed-matter physics. In this review we address the collective behavior of systems of multicomponent scalar fields interacting with gauge fields, which can be already present in the underlying microscopic system or emerge only at criticality. The interplay between local gauge and global symmetries determines the phase diagram, the nature of the Higgs phases, and the nature of phase transitions between the high-temperature disordered and the low-temperature Higgs phases. However, additional crucial features determine the universal properties of the critical behavior at continuous transitions. Specifically, their nature also depends on the role played by the gauge modes at criticality. Effective (Abelian or non-Abelian) gauge Higgs field theories emerge when gauge modes develop critical correlations. On the other hand, a more standard critical behavior, which admits an effective description in terms of Landau-Ginzburg-Wilson $\Phi^{4}$ theories, occurs when gauge-field modes are short ranged at the transition. In the latter case, gauge fields only prevent non-gauge invariant correlation functions from becoming critical. This review covers the recent progress made in the study of Higgs systems with Abelian and non-Abelian gauge fields. We discuss the equilibrium thermodynamic properties of systems with a classical partition function, focusing mainly on three-dimensional systems, and only briefly discussing two-dimensional models. However, by using the quantum-to-classical mapping, the results on the critical behavior for classical systems in $D=d+1$ dimensions can be extended to quantum transitions in $d$ dimensions.
@@
Abelian and non-Abelian gauge theories with scalar fields, Higgs mechanism, Phase transitions, Critical behavior in the presence of gauge symmetries, Landau-Ginzburg-Wilson paradigm, Charged transitions with critical gauge fields, Topological transitions.
@@ 1.1 Introduction
Many collective phenomena in condensed-matter physics are modelled by using effective Abelian Higgs (AH) theories, which describe a $d$-dimensional system of degenerate $N$-component scalar fields minimally coupled with an Abelian U(1) gauge field. We mention transitions in superconductors, see, e.g., Refs. [18, 19], and in quantum SU($N$) antiferromagnets, see, e.g., Refs. [27, 28, 29, 30, 31, 32, 20], and some unconventional quantum transitions characterized by the so-called deconfined quantum criticality, see, e.g., Refs. [33, 34, 35, 36, 37, 38, 39, 40]. The phase structure and universal features of AH models have been extensively studied, see, e.g., Refs. [18, 19, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 35, 76, 77, 78, 30, 79, 31, 80, 32, 81, 82, 83, 84, 20, 85, 86, 87, 88, 89, 90, 91, 21, 92, 93, 94], paying particular attention to the role of the gauge fields and of the related topological features, like monopoles and Berry phases, which cannot be captured by effective Landau-Ginzburg-Wilson (LGW) theories [5, 95, 6, 2, 96] with gauge-invariant scalar order parameters, see, e.g., Refs. [33, 20, 21].
@@
From the theoretical point of view, the existence of continuous phase transitions that cannot be described by the standard gauge-invariant LGW paradigm is related to the nonperturbative infrared behavior of strongly coupled gauge theories. In particular, the existence of infrared stable charged fixed points (FPs) of the renormalization-group (RG) flow implies that gauge-field correlations can also be critical at some continuous transition. At a fundamental level, the existence of these nontrivial transitions also provides a way to define gauge field theories in a nonperturbative framework.
Over the past decade, researchers with different backgrounds, working in high-energy, statistical, and condensed-matter physics, have made important steps forward in the theoretical understanding of the nonperturbative physics of Abelian and non-Abelian gauge Higgs models in less than four dimensions. In light of this steady progress, a survey of the field seems to be useful for two different reasons: First, one can present the state of the art in the field; second, there is the possibility to collect and discuss in a unified fashion results that appeared in different contexts. This will also promote a more efficient communication and interdisciplinary interactions among researchers working in different areas.
To the best of our knowledge, a review on lower-dimensional, i.e., three-dimensional (3D) and two-dimensional (2D), AH and NAH theories is still missing, and we would like to fill such gap. Of course, some of the aspects we plan to cover have already been reviewed— emerging gauge fields and deconfined criticality are obvious examples, see, e.g., Refs. [20, 21]. However, the existing reviews focus on specific issues and do not report a general discussion of lower-dimensional lattice AH and NAH theories. In this review we thoroughly examine the phenomenology of AH and NAH systems, in particular their phase diagram and the nature of the critical transitions observed when varying the gauge and global symmetries.
@@ 1.2 Plan
This review focuses on lower-dimensional AH and NAH theories, in which multicomponent scalar fields are coupled with Abelian and non-Abelian gauge fields. We mostly discuss these models in a statistical lattice gauge framework [101, 2, 10, 3, 4], in which the partition function is defined as a sum over classical lattice configurations, thus providing a regularization of the functional path integral and, possibly, paving the way for a nonperturbative definition of the corresponding statistical field theory. The quantum-to-classical mapping allows us to apply the classical results for the critical behavior (equivalently, continuum limit) in $(d+1)$ dimensions to quantum transitions of many-body systems in $d$ dimensions, see, e.g., Refs. [14, 102]. In particular, since we consider isotropic $(d+1)$-dimensional statistical models, the results can be applied to the large class of continuous quantum transitions characterized by a dynamic exponent $z=1$, i.e., to transitions in which the critical gap decays as the inverse of the spatial size [14, 102]. We mainly consider 3D classical systems, but we also briefly discuss 2D systems.
@@
The review covers the recent developments in the studies of lattice AH and NAH systems (LAH and LNAH, respectively). We discuss their phase diagram, the main features of their Higgs ordered phases [103, 104, 105, 106], where gauge correlations are gapped and nonlocal charged operators may condense, and the nature of their phase transitions, paying particular attention to transitions characterized by a nontrivial interplay of gauge and scalar modes. We will show that the standard effective LGW $\Phi^{4}$ field theory approach [95, 2, 96], which is commonly applied to critical phenomena in statistical systems without gauge symmetries, is not always able to describe transitions in gauge models. In some cases different effective approaches are needed to correctly capture the nature of the critical behavior. We pay particular attention to the comparison with the predictions of the corresponding AH and NAH field theories (AHFTs and NAHFTs, respectively), to understand when, and how, the critical behavior of lattice systems is effectively described by a corresponding field-theory model. If this is the case, the lattice model provides a way, the so-called continuum limit, to define nonperturbatively the field theory.
@@
In Secs. 2 and 3 we begin by discussing the possible types of critical behavior that emerge at continuous transitions in the presence of gauge symmetries. This requires an extension of the standard LGW paradigm, in particular when gauge modes play an active role at criticality. Secs. 4-6 focus on systems characterized by an Abelian U(1) gauge invariance, i.e., the AHFTs and the corresponding LAH models. We consider models with both noncompact and compact formulations of the gauge fields, which turn out to present significantly different features. Lattice models with discrete Abelian gauge symmetries are addressed in Sec. 7. Field theories and lattice systems with non-Abelian gauge symmetries are discussed in Secs. 8 and 9, respectively. Two-dimensional models, with Abelian and non-Abelian gauge symmetries, are discussed in Sec. 10. Finally, in Sec. 11 we discuss the effects of perturbations explicitly breaking the gauge symmetry, an issue which is particularly relevant for systems in which the gauge symmetry effectively emerges at the transition but which are not gauge invariant at the microscopic level. A more detailed description of the content of the review follows.
@@
    In Sec. 3, we discuss the key features of the critical behavior of statistical systems with local gauge symmetries. We argue that different types of critical behaviors may be realized. There are transitions driven by scalar fields whose nature depends on the role played by the gauge modes, in particular whether they develop or not critical correlations, and transitions driven by topological gauge modes without local order parameters. The effective description of these different critical behaviors requires different effective approaches, such as the standard gauge-invariant LGW approach when gauge fields are not critical, or the more complex gauge field theory (GFT) approach with explicit gauge fields in the presence of critical gauge correlations. This distinction is crucial to understand which statistical field theory (SFT) is actually realized by the various continuous phase transitions occurring in the presence of gauge symmetries.
@@
    In Sec. 4 we focus on AHFTs, or scalar quantum electrodynamics (QED), which describe $d$-dimensional systems of degenerate $N$-component scalar fields minimally coupled with Abelian U(1) gauge fields. In particular, we consider systems with global SU($N$) and SO($N$) symmetry. We review the main features of their RG flow, whose stable charged FP is expected to describe the critical behavior at charged continuous transitions, when both scalar and gauge fields are critical.
@@
    In Sec. 5 we analyze lattice models corresponding to the AHFTs discussed in Sec. 4, which are characterized by the same global and local symmetries. We consider 3D LAH models in which a noncompact Abelian gauge field is coupled with an $N$-component scalar field, providing a straightforward lattice discretization of the AHFT. We review results for their phase diagrams, which present various phases, including a Higgs phase, and for their critical behavior along the various transition lines. Different critical behaviors occur, which can be described by using the different approaches outlined in Sec. 3. In particular, the transitions separating the Coulomb and Higgs phases realize the critical continuum limit associated with the RG flow of the AHFTs for a sufficiently large number of scalar components.
@@
    In Sec. 8 we review NAHFTs, in which multiflavor scalar fields are coupled with non-Abelian gauge fields. We discuss models with SU($N_{c}$) and SO($N_{c}$) gauge symmetries, with scalar fields transforming in different representations of the gauge group, such as the fundamental and adjoint representations. We outline the known features of their RG flow, whose stable infrared FPs are expected to describe the critical behavior at some continuous phase transitions in the corresponding lattice statistical models.
@@
    In Sec. 9 we consider LNAH models, which are lattice non-Abelian gauge theories with multicomponent degenerate scalar fields. We present results for their phase diagram, paying particular attention to their low-temperature Higgs phases. We discuss the interplay between non-Abelian gauge symmetries and global symmetries, which crucially determines the properties of the various Higgs phases, and therefore the nature of the phase transitions between the disordered and the Higgs phases. In particular we consider SU($N_{c}$) gauge theories with multicomponent scalar fields in the fundamental and adjoint representations of the gauge group, two choices which lead to different Higgs phases. Some of the transitions between the disordered and the Higgs phases are continuous and can be associated with the charged FP of the corresponding NAHFT. We close this section by briefly discussing SO($N_{c}$) gauge theories with multiflavor scalar matter.
@@
    In Sec. 10 we review results for the critical continuum limit of 2D LAH and LNAH models with continuous global symmetries. They show that also the critical continuum limit of 2D LAH and LNAH models in the zero-temperature limit arises from a nontrivial interplay between gauge and global symmetries. The emerging results lead us to conjecture that the zero-temperature critical behavior, and therefore the continuum limit, of 2D lattice gauge models with scalar fields belongs to the zero-temperature universality classes associated with 2D field theories ($\sigma$ models) defined on symmetric spaces.
@@
    In Sec. 11 we discuss the effects of perturbations that give rise to an explicit gauge-symmetry breaking (GSB) in lattice gauge theories. This issue is discussed in LAH models with noncompact and compact gauge fields in the presence of photon-mass terms. We argue that their effect is irrelevant at the continuous transitions where gauge correlations are not critical, and the critical behavior can be described by effective LGW theories with gauge-invariant order-parameter fields. On the other hand, photon-mass terms are expected to be relevant at charged transitions where the gauge modes develop critical correlations.
@@
    In A we report an overview of the homogeneous scaling laws expected at generic continuous phase transitions, as inferred from the RG theory of critical phenomena. We report the RG scaling relations valid in the thermodynamic limit and in the finite-size scaling (FSS) limit, at critical and multicritical points. We also briefly discuss FSS at first-order transitions.
@@
    For reference, in B we report the presently most accurate estimates of the critical exponents of the 3D Ising, XY, and O(3) vector (Heisenberg) universality classes, which are often referred to in this review.
@@
    In C we present some details on the mean-field analyses of the Higgs phases of LNAH models with SU($N_{c}$) gauge symmetry and $N_{f}$ flavors in the fundamental representation.
@@
Another important issue that is not addressed in this review concerns the new features that may arise when fermionic fields are added to lattice gauge Higgs models. This topic is quite interesting, since models with emerging gauge symmetries in condensed-matter physics often involve fermionic excitations beside scalar modes. We believe that this issue may provide further interesting developments, that would extend our present understanding of the phenomenology of the transitions in the presence of gauge symmetries. Recent developments, mostly in field-theoretical frameworks, have been already reported in the literature; e.g., in the review [20] and, in particular, in the recent book [16], which also discusses boson-fermion and fermion-fermion dualities. However, 3D lattice gauge Higgs models including fermionic fields have been much less studied so far. Therefore, we prefer to restrict our review to lattice gauge models with scalar matter only, for which there is already a large amount of results that have not been surveyed yet.
To conclude, we may summarize the content of this review as follows. It is meant to provide a comprehensive presentation of the state of the art of the research on the nonperturbative behavior of low-dimensional (2D and 3D) Abelian and non-Abelian Higgs theories. For this purpose we present a survey of results for gauge Higgs theories, focusing mainly on 2D and 3D lattice realizations, in which the partition function is defined as a sum over classical lattice configurations. These models are particularly effective to understand the nonperturbative properties of gauge-Higgs theories and the universal features of their transitions. We emphasize that their critical behaviors allow one to obtain nonperturbative definitions of the statistical field theories sharing the same global and local gauge symmetries, which can also be investigated by standard field-theoretical approaches.
In the following we report a list of the abbreviations and symbols used throughout the review (in alphabetic order):